### v1

# Deterministic Self-Stabilising Leader Election for Programmable Matter with Constant Memory

Abstract
The problem of electing a unique leader is central to all distributed systems, including programmable matter systems where particles have constant size memory. In this paper, we present a silent self-stabilising, deterministic, stationary, election algorithm for particles having constant memory, assuming that the system is simply connected. Our algorithm is elegant and simple, and requires constant memory per particle. We prove that our algorithm always stabilises to a configuration with a unique leader, under a daemon satisfying some fairness guarantees (Gouda fairness [26]). We use the special geometric properties of programmable matter in 2D triangular grids to obtain the first self-stabilising algorithm for such systems. This result is surprising since it is known that silent self-stabilising algorithms for election in general distributed networks require $\Omega(\log{n})$ bits of memory per node, even for ring topologies [19].
keywords:
Leader Election, Programmable Matter, Self-Stabilisation, Silent, Deterministic, Unique Leader, Simply Connected, Gouda fair Daemon, Constant Memory
1 Introduction
Leader election (introduced by Le Lann [31]) allows to distinguish a unique process in the system as a leader. The leader process can then act as an initiator or a coordinator, for solving other distributed problems. Thus, election algorithms are often used as building blocks for many problems in this domain. We are interested in deterministic election algorithms that are self-stabilising. Since the seminal work of Dijkstra [17], the self-stabilisation paradigm has been thoroughly investigated (see [18] for a survey). A distributed algorithm is self-stabilising if when executed on a distributed system in an arbitrary global initial configuration, the system eventually reaches a legitimate configuration. Self-stabilising protocols are able to autonomously recover from transient memory failures, without external intervention. A self-stabilising algorithm is silent if the system always reaches a configuration where the processes no longer change their states. In the self-stabilising setting, leader election is particularly important, as many self-stabilising algorithms rely on the existence of a distinguished node.
The concept of silent self-stabilising algorithms is also related to proof-labelling schemes [30] where each node is given a local certificate to verify certain global properties of the system (e.g., the existence of a unique leader). Each node can check its own certificate and those of its neighbours to verify it is in a correct configuration. If the global configuration is incorrect, at least one node should be able to detect an inconsistency using the local certificates. In this case, this node will change its state, leading its neighbours to change their states and so on, until the system stabilises to a correct configuration. Blin et al. [5] proved that from any proof-labelling scheme where each process has a certificate of size $\ell$, one can build a silent self-stabilising algorithm using $O(\ell+\log n)$ bits of memory per process, where $n$ is the number of processes in the network. One of the standard techniques for self-stabilising leader election is to build a spanning tree rooted at the leader, with all other nodes pointing towards their parent in the tree. In order to detect cycles when the system is in an incorrect state, the local certificate at each node includes the hop-distance to the root, in addition to the pointer to the parent. So the size of the certificate depends on the size of the system.
In this paper, we consider programmable matter systems which are distributed systems consisting of small, intelligent particles that connect to other particles to construct a given shape and can autonomously change shapes according to input signals. Such systems should be scalable to arbitrary sizes, so the particles are assumed to have constant size memory independent of the size of the system, similar to finite state automata. This requirement also implies that the particles are anonymous (i.e., do not have unique identifiers) and all communication is limited to $O(1)$ size messages. One well-studied model for programmable matter is the Amoebot model [14] where particles operate on a triangular grid (see Section 1.2 for details). Leader election is a well studied problem in this model. When the system is simply connected, there are stationary deterministic algorithms for election based on the erosion approach [16] where the algorithm starts by deactivating particles on the boundary and moving inwards, until the last active node becomes the leader. This approach works under the minimum assumptions on the system and is the inspiration for our work.
Tolerating faults is important for programmable matter, however none of the existing algorithms for election in these systems are self-stabilising. The question is: given the constant memory of particles, is it still possible to obtain a self-stabilising algorithm for programmable matter, using other properties of such systems? We answer this question in the affirmative, at least for simply connected programmable matter systems, showing that in this case, a deterministic silent self-stabilising algorithm for leader election is indeed possible. We use the property of such systems that there is a unique boundary in the system that is well defined, such that any particle can determine whether it is on the boundary.
1.1 Our results
We present a silent self-stabilising, deterministic, stationary, election algorithm for constant-memory particles, in a simply connected system. We prove our algorithm always stabilises to a unique leader configuration, under a scheduler with some fairness guarantees.
We first present a proof labelling scheme ensuring the existence of a unique leader. Our certificate orients the edges of the network and a configuration is valid when: every edge is oriented, every particle has at most three outgoing edges, outgoing edges appear consecutively around each particle and there are no directed triangles. Note that our certificate does not ensure that the global orientation of the network is acyclic. However, using the geometric properties of the configuration, we are able to show that any valid configuration has a unique sink. As we are interested in a constant memory algorithm, one cannot transform our proof labelling scheme into a self-stabilising algorithm using [5]. However, we design a very simple algorithm to orient the edges of the system. We show that under our fairness assumption, one always reaches a valid configuration and that this configuration contains a unique sink that is designated as the leader.
Following the classification of [20], our scheduler is Gouda fair [26]: for any configuration $C$ that appears infinitely often in the execution, any successor $C^{\prime}$ of $C$ also appears infinitely often. Since each particle has constant memory, there exists only a finite number of global configurations of the system. In this setting, Gouda fairness ensures that any configuration that is infinitely often reachable is eventually reached. Observe that a scheduler that at each step activates a particle chosen uniformly at random is a Gouda fair sequential scheduler.
We do not assume that there exists an agreement on the orientation of the grid, or even on its chirality. Observe that without simple connectivity and without agreement on orientation or chirality, it is possible to construct arbitrarily large rings of even size where all processes have the same geometric information about the system (see Fig. 1). In this setting, the results of [19] show that there is no silent self-stabilising leader election algorithm using constant memory. This is part of our motivation for considering simply connected systems.
Figure 1: An 18-particle ring where for each particle, the occupied neighbours are reached through port numbers 2 and 4. Two nodes have the same colour if they agree on the grid orientation.
1.2 Related Work
In general networks, there is no self-stabilising leader election algorithm where each process has a constant memory. More precisely, Dolev et al. [19] established that any silent self-stabilising algorithm electing a leader in the class of rings requires $\Omega(\log n)$ bits of memory per process (where $n$ is the size of the ring). This lower bound only uses the assumption that there exists a silent correct configuration and holds for any kind of scheduler. More recently, Blin et al. [4] showed that non-silent self-stabilising algorithms require $\Omega(\log\log n)$ bits of memory per process in order to elect a leader synchronously in the class of rings. Note that these lower bounds are tight in the sense that for rings, there exist silent (resp. non-silent) self-stabilising leader election algorithms using $O(\log n)$ (resp., $O(\log\log n)$) bits of memory per process [11, 6]. Constant memory self stabilising algorithms for rings can be designed under special assumptions, as in [29] which gives an algorithm for prime sized rings assuming a sequential scheduler. However, [19] established that this algorithm cannot be made silent.
There exists a large literature about distributed systems where each process has finite memory. Cellular automata, introduced in the 40s in [33] are one of the best known models of this kind. More recently, numerous papers have been devoted to population protocols introduced in [1]. In this model, there is a population of finite-state agents and at each step, a scheduler picks two agents that jointly update their states according to their current states. The scheduler satisfies the same fairness condition as the one we consider in this paper: any configuration that is infinitely often reachable is eventually reached. In this setting, there exist election protocols using only two states when all agents start in the same state. However, when considering self-stabilising leader election in this setting, Cai et al. [9] showed that a protocol using $n-1$ states cannot solve the problem in a population of $n$ agents. This shows that even with a Gouda fair scheduler, it is not always possible to solve the leader election problem in a self-stabilising way when processes have constant memory.
Programmable matter was introduced in [32] and has since gained popularity. Several models have been introduced, such as [27, 34, 23]. In this paper we consider the well studied Amoebot model [14, 12]. In this model, constant-memory computational entities, called particles, operate in a triangular grid. Each node of the grid is occupied by at most one particle and particles can determine whether nodes at distance one are occupied by particles. Each particle can communicate with its neighbours by reading their respective registers. It is usually assumed that particles do not have any global sense of direction, while some papers (unlike this paper) assume that the particles have a common sense of rotational orientation, called chirality (e.g., [21, 3]) or that particles agree on a common direction (e.g., [10]). Particles in the Amoebot model have the ability to move to neighbouring nodes (e.g., [21, 22]), which we do not use here. The problem of Leader Election has been studied in the specific context of programmable matter in both two dimensional (e.g., [3, 16]) and three dimensional settings (e.g., [25, 8]) and both deterministic (e.g., [21]) and randomized algorithms (e.g., [15]) have been proposed. The existing algorithms for leader election in programmable matter can be categorized based on the use of two main techniques: erosion (e.g., [16, 24]) and message passing on boundaries (e.g., [3, 15]).
Research on self-stabilisation in the programmable matter setting is more limited. In [15], a randomised leader election algorithm is given and the authors discuss the possibility of making it self-stabilising by combining it with techniques from [2, 28]. However, it is assumed that particles have $O(\log^{*}n)$ memory. In the same paper, it is also argued that self-stabilisation in programmable matter is not possible for problems where movement is needed, as the system can become permanently disconnected. More recently, in [13] a self-stabilising algorithm for constructing a spanning forest was introduced. In this case, the algorithm is deterministic and particles have constant memory. However, it is assumed that at least one non-faulty special particle always remains in the system. The need to extend the Amoebot model to also address self-stabilising algorithms is also discussed in [12].
2 Model
Let $G_{\Delta}$ be an infinite regular triangular grid where each node has six neighbours. A connected particle system, $\mathcal{P}$, is simply connected if $G_{\Delta}\backslash\mathcal{P}$ is connected. We assume each node of the simply connected $\mathcal{P}$ contains exactly one particle. We call nodes that are in $\mathcal{P}$ occupied and those that are not in $\mathcal{P}$, empty. Each particle is anonymous, has constant memory and is stationary (i.e., does not move). A particle is incident to six ports, leading to consecutive neighbouring nodes in $G_{\Delta}$. Each port is associated with a label so that ports $i$ and $i+1\mod 6$ lead to neighbouring nodes. A particle knows if each port leads to an occupied or empty node. For each occupied neighbour $q$, particle $p$ knows the label assigned by $q$ to the edge $qp$. Each particle has a constant-size register with arbitrary initial contents. A particle can read the register of each occupied neighbour but can only write in its own register. All particles are inactive unless activated by the scheduler. An activated particle reads the contents of its register and the register of each of the neighbouring particles. Based on this information it updates the contents of its own register according to the given algorithm.
We call $\mathcal{P}$, the support of the particle system. The configuration $C$ of the system at any time, consists of the set $\mathcal{P}$ and the contents of the registers of each particle in $\mathcal{P}$. A distributed algorithm $\mathcal{A}$ is a set of local rules that particles execute. The rules of the algorithm depend only on the content of the registers of the particles and of its neighbours and they modify only the register of the particle. For an algorithm $\mathcal{A}$, a configuration $C$, and a particle $p$, we say that $p$ is activable in $C$, if the execution of $\mathcal{A}$ modifies the contents of the register of $p$. For two configurations, $C$ and $C^{\prime}$ that have the same support, we say that $C^{\prime}$ is a successor (resp. sequential successor) of $C$ if there exists a non-empty set of activable particles $Q$ (resp. an activable particle $p$) in $C$ such that, when all particles in $Q$ execute (resp. when $p$ executes) $\mathcal{A}$, $C^{\prime}$ is obtained. An execution $\mathcal{S}$ is an infinite sequence of configurations $\mathcal{S}=C_{0},C_{1},\ldots$ such that for any $i$, $C_{i}$ and $C_{i+1}$ have the same support and either there exists a non-empty set $Q_{i}$ of activable particles such that when all particles from $Q_{i}$ execute $\mathcal{A}$ concurrently in $C_{i}$, $C_{i+1}$ is obtained, or there is no activable particle and $C_{i+1}=C_{i}$. If $|Q_{i}|=1$ for any $i$, then we say that $\mathcal{S}$ is a sequential execution. If there exists a step where $C_{i+1}=C_{i}$, we call $C_{i}$ a final configuration. An execution is Gouda fair [20, 26] if for any configuration $C$ that appears infinitely often in the execution, any successor $C^{\prime}$ of $C$ also appears infinitely often. A sequential execution is sequential Gouda fair if for any configuration $C$ that appears infinitely often in the execution, any successor $C^{\prime}$ of $C$ that is obtained by activating one particle of $C$ also appears infinitely often. The notion of valid configurations depends on the algorithm. In the next section, we define the valid configurations we consider in this paper. An algorithm $\mathcal{A}$ is silent if the set of valid configurations is precisely the set of final configurations.
Remark 2.1.
In our analysis we show that our algorithm is correct under a sequential Gouda fair scheduler but from [26, Theorem 8] our result also holds under a Gouda fair scheduler (that is not necessarily sequential).
We now present some notations and observations about the geometry of the system. Let $v$ and $v^{\prime}$ be two neighbouring nodes in $\mathcal{P}$. We say that an edge that is oriented from $v$ to a neighbouring node $v^{\prime}$ is outgoing for $v$ and incoming for $v^{\prime}$. We write $\overrightarrow{vv^{\prime}}$ to denote an edge directed from $v$ to $v^{\prime}$ and $vv^{\prime}$ to denote an undirected edge or an edge whose orientation is not known. Particles with at least one neighbour that is not in $\mathcal{P}$ are on the boundary. Since $\mathcal{P}$ is simply connected, there exists only one boundary in the system. Let $p$ be a particle on the boundary. We say that $p$ is pending if $p$ has a unique neighbouring particle in $\mathcal{P}$. We say that $p$ is an articulation point if the removal of $p$ disconnects $\mathcal{P}$. If $p$ is neither pending, nor an articulation point, then $p$ is incident to two distinct edges $pq$, $pr$ on the boundary of $\mathcal{P}$. In this case, since $\mathcal{P}$ is simply connected, there is a path of particles in the 1-neighbourhood of $p$ from $q$ to $r$. We say that $p$ is on a $\theta\in\{60^{\circ},120^{\circ},180^{\circ},240^{\circ}\}$ angle to denote the angle that is formed when moving from $q$ to $r$ around $p$ and no empty nodes are encountered. By slight abuse of notation, we also call a particle on a $\theta$ angle a $\theta$ particle.
It is easy to see that a particle on the boundary cannot be on a $300^{\circ}$ angle, otherwise $q$ and $r$ are adjacent and $p$ is not on the boundary, a contradiction. $\mathcal{P}$ is 2–connected if it does not contain any articulation point. Notice that in systems with at least three particles, a system with no articulation point does not contain any pending particle. In a 2–connected particle system, the following observation implies that there should be a $60^{\circ}$ or a $120^{\circ}$ particle.
{observation}
If $\mathcal{P}$ is 2–connected and contains at least three particles, particles on the boundary satisfy the formula $62n_{60}+n_{120}-n_{240}=$ 6, where $n_{\theta}$ is the number of $\theta$ particles on the boundary.
Proof 2.2.
If $\mathcal{P}$ is 2–connected, the boundary forms a simple polygon. We know that the sum of internal angles of a simple polygon is $(n-2)\pi$, where $n$ is the number of vertices of the polygon. So $(n_{60}+n_{120}+n_{180}+n_{240}-2)\pi=n_{60}\frac{\pi}{3}+n_{120}\frac{2\pi}{3}+n_{180}\pi+n_{240}\frac{4\pi}{3}$, that is, $62n_{60}+n_{120}-n_{240}=$ 6.
Lemma 2.3.
In any simply connected particle system $\mathcal{P}$ with at least two particles, the boundary of $\mathcal{P}$ contains one of the following:
  1. 1.
    a pending particle, or
  2. 2.
    a $60^{\circ}$ particle, or
  3. 3.
    two $120^{\circ}$ particles that are connected by a path of $180^{\circ}$ particles on the boundary.
Proof 2.4.
A block is a 2–connected component of $\mathcal{P}$. As $\mathcal{P}$ contains at least two particles, each block is either an edge or it contains at least three particles. The block tree of $\mathcal{P}$ is a tree where each vertex is a block and there is an edge between two blocks if they share a vertex (i.e., an articulation point of $\mathcal{P}$). A leaf, $\mathcal{P}^{\prime}$, of the block tree is a 2-connected component of $\mathcal{P}$ and contains a unique articulation point $p^{\prime}$ of $\mathcal{P}$. If $\mathcal{P}^{\prime}$ contains precisely two particles $p^{\prime}$ and $q^{\prime}$, then $p^{\prime}$ is the unique neighbour of $q^{\prime}$ in $\mathcal{P}$ and $q^{\prime}$ is a pending particle, as in Case 1.
Suppose $\mathcal{P}^{\prime}$ contains at least three particles. Since $\mathcal{P}^{\prime}$ is 2-connected, every particle on the boundary of $\mathcal{P}^{\prime}$ is a $\theta\in\{60^{\circ},120^{\circ},180^{\circ},240^{\circ}\}$ particle. Any $\theta$ particle $p\neq p^{\prime}$ of $\mathcal{P}^{\prime}$ is also a $\theta$ particle of $\mathcal{P}$. So a $60^{\circ}$ particle $p\neq p^{\prime}$ in $\mathcal{P}^{\prime}$, is a $60^{\circ}$ particle in $\mathcal{P}$, which is Case 2.
Suppose now that in $\mathcal{P}^{\prime}$, any boundary particle $p$ different from $p^{\prime}$ is a $\theta$ particle with $\theta\in\{120^{\circ},180^{\circ},240^{\circ}\}$. Let $n^{\prime}_{120}$, $n^{\prime}_{180}$, $n^{\prime}_{240}$ be respectively the number of $120^{\circ}$, $180^{\circ}$, $240^{\circ}$ particles in $\mathcal{P}^{\prime}$ that are different from $p^{\prime}$. Since $p^{\prime}$ is an articulation point, $p^{\prime}$ cannot have more than three consecutive particle neighbours. Consequently, in $\mathcal{P}^{\prime}$, $p^{\prime}$ is either a $60^{\circ}$ or a $120^{\circ}$ particle. If $p^{\prime}$ is a $60^{\circ}$ particle in $\mathcal{P}^{\prime}$, from Observation 2, we have $62+n^{\prime}_{120}-n^{\prime}_{240}=$ 6. If $p^{\prime}$ is a $120^{\circ}$ particle in $\mathcal{P}^{\prime}$, from Observation 2, we have $n^{\prime}_{120}+1-n^{\prime}_{240}=6$. In both cases, we then have $n^{\prime}_{120}\geq n^{\prime}_{240}+4$. Let $p_{1},\ldots,p_{n^{\prime}_{120}}$ be the $120^{\circ}$ particles of $\mathcal{P}^{\prime}$ in the order in which they appear when we move on the boundary of $\mathcal{P}^{\prime}$ starting from $p^{\prime}$ (i.e., $p^{\prime}$ appears between $p_{n^{\prime}_{120}}$ and $p_{1}$). Since $n^{\prime}_{120}\geq n^{\prime}_{240}+4>n^{\prime}_{240}+1$, there exists an index $11\leq i\leq n^{\prime}_{120}-$ 1 such that only $180^{\circ}$ particles appear on the boundary of $\mathcal{P}^{\prime}$ between $p_{i}$ and $p_{i+1}$. Since all these $180^{\circ}$ particles are also $180^{\circ}$ particles on the boundary of $\mathcal{P}$, we are in Case 3.
We now explain how two adjacent particles in a triangle can detect each other’s chirality. The label $\lambda(\Pi)$ of a path $\Pi=(p_{1},p_{2},\ldots,p_{k})$ in the graph induced by the particles is a sequence of pairs of labels $(\mathsf{a}_{1},\mathsf{b}_{2}),(\mathsf{a}_{2},\mathsf{b}_{3}),\ldots,(\mathsf{a}_{k-1},\mathsf{b}_{k})$ where for each $i$, $\mathsf{a}_{i}$ (resp. $\mathsf{b}_{i}$) is the port connecting $p_{i}$ to $p_{i+1}$ (resp. $p_{i-1}$). Following the definitions of [35], we define $\texttt{view}_{k}(p)$, to be the set of labels $\lambda(\Pi)$ of paths $\Pi$ starting at $p$ of length at most $k$. Note that for each $1\leq j\leq k$, if both $(\mathsf{a}_{1},\mathsf{b}_{2}),(\mathsf{a}_{2},\mathsf{b}_{3}),\ldots,(\mathsf{a}_{j},\mathsf{b}_{j+1})$ and $(\mathsf{a}_{1},\mathsf{b}_{2}),(\mathsf{a}_{2},\mathsf{b}_{3}),\ldots,(\mathsf{a}_{j},\mathsf{b}_{j+1}^{\prime})$ belong to $\texttt{view}_{k}(p)$, then $\mathsf{b}_{j+1}=\mathsf{b}_{j+1}^{\prime}$. From [7], for any constant $k$, each particle $p$ can construct $\texttt{view}_{k}(p)$ in a self stabilising way with constant memory.
Lemma 2.5.
For any triangle of particles $pqr$, $p$ can infer the chirality of $q$ from $\emph{view}_{3}(p)$.
Proof 2.6.
In the following, for a particle $p$, we let $\{\mathsf{p}_{i}\mid 0\leq i\leq 5\}$ be the set of ports incident to $p$ and we assume that either $\mathsf{p}_{i+1}=\mathsf{p}_{i}+1$ for each $0\leq i\leq 5$, or $\mathsf{p}_{i+1}=\mathsf{p}_{i}-1$ for each $0\leq i\leq 5$ (where additions are made modulo $6$). We will use the following observation.
Claim 1.
If $pqr$ is a triangle, then the ports connecting $r$ to $p$ and $q$ are consecutive.
Consider a triangle $pqr$. Let $\mathsf{p}_{1}$ (resp. $\mathsf{q}_{1}$) be the port through which $p$ (resp. $q$) is connected to $q$ (resp. $p$). Further, let $p$ (resp. $r$) be connected to $r$ (resp. $p$) through $\mathsf{p}_{0}$ (resp. $\mathsf{r}_{1}$). Observe that if $p$ learns the port through which $q$ is connected to $r$, it also learns the chirality of $q$. Note that by Claim 1, this port is either $\mathsf{q}_{0}$ or $\mathsf{q}_{2}$ and the port from $r$ to $q$ is either $\mathsf{r}_{0}$ or $\mathsf{r}_{2}$. Notice that if $r$ is the only common neighbour of $p$ and $q$, then only one of $\{(\mathsf{p}_{1},\mathsf{q}_{1}),(\mathsf{q}_{0},\mathsf{x})\mid 0\leq\mathsf{x}\leq 5\}\cup\{(\mathsf{p}_{1},\mathsf{q}_{1}),(\mathsf{q}_{2},\mathsf{x})\mid 0\leq\mathsf{x}\leq 5\}$ is in $\texttt{view}_{3}(p)$ and $p$ can then infer the chirality of $q$. Suppose now that $p$ and $q$ have two common neighbours $r$ and $r^{\prime}$.
Claim 2.
The edge $qr$ is labelled $(\mathsf{q}_{2},\mathsf{r}_{0})$ if and only if the following formula holds:
$(\mathsf{p}_{1},\mathsf{q}_{1}),(\mathsf{q}_{2},\mathsf{r}_{0}),(\mathsf{r}_{1},\mathsf{p}_{0})\in\texttt{view}(p)\text{ }\land$
$(\mathsf{p}_{0},\mathsf{r}_{1}),(\mathsf{r}_{0},\mathsf{q}_{2}),(\mathsf{q}_{1},\mathsf{p}_{1})\in\texttt{view}(p)\text{ }\land$
$\Bigl{[}(\mathsf{p}_{1},\mathsf{q}_{1}),(\mathsf{q}_{0},\mathsf{r}_{2})\notin\texttt{view}(p)\text{ }\lor$
$(\mathsf{p}_{2},\mathsf{r}_{5})\notin\texttt{view}(p)\text{ }\Bigr{]}$
{claimproof}
Notice that the configuration is either the one on Figure 2(a) or on Figure 2(b), since $p$ knows the labels of its incident edges.
$p$$\mathsf{p}_{1}$$\mathsf{p}_{0}$$\mathsf{p}_{2}$$r$$\mathsf{r}_{1}$$q$$\mathsf{q}_{1}$$\mathsf{q}_{2}$$r^{\prime}$
$p$$\mathsf{p}_{1}$$\mathsf{p}_{0}$$p_{2}$$r$$\mathsf{r}_{1}$$q$$\mathsf{q}_{1}$$\mathsf{q}_{2}$$r^{\prime}$
Figure 2: The 1-view of $p$ and the two possible orientations of $q$.
First let us suppose that the edge $qr$ is labelled $(\mathsf{q}_{2},\mathsf{r}_{0})$. Then, the first two expressions of the formula are satisfied. Let us suppose $(\mathsf{p}_{1},\mathsf{q}_{1}),(\mathsf{q}_{0},\mathsf{r}_{2})\in\texttt{view}_{3}(p)$. Then from Claim 1, $qr^{\prime}$ is labelled $(\mathsf{q}_{0},\mathsf{r}_{2})$ and $pr^{\prime}$ is either labelled $(\mathsf{p}_{2},\mathsf{r}_{1})$ or $(\mathsf{p}_{2},\mathsf{r}_{3})$. In either case, $(\mathsf{p}_{2},\mathsf{r}_{5})\notin\texttt{view}_{3}(p)$ and the formula is satisfied.
Let us now suppose that the formula is satisfied and assume that $qr$ is not labelled $(\mathsf{q}_{2},\mathsf{r}_{0})$. Then by Claim 1, $qr$ is labelled either $(\mathsf{q}_{2},\mathsf{r}_{2})$, or $(\mathsf{q}_{0},\mathsf{r}_{0})$, or $(\mathsf{q}_{0},\mathsf{r}_{2})$. Note that the first two cases are impossible since $(\mathsf{p}_{1},\mathsf{q}_{1}),(\mathsf{q}_{2},\mathsf{r}_{0})$ and $(\mathsf{p}_{0},\mathsf{r}_{1}),(\mathsf{r}_{0},\mathsf{q}_{2})$ belong to $\texttt{view}_{3}(p)$. Consequently, $qr$ is labelled $(\mathsf{q}_{0},\mathsf{r}_{2})$ and since $(\mathsf{p}_{1},\mathsf{q}_{1})(\mathsf{q}_{2},\mathsf{r}_{0})\in\texttt{view}_{3}(p)$, by Claim 1, $qr^{\prime}$ is labelled $(\mathsf{q}_{2},\mathsf{r}_{0})$, and the label of $pr^{\prime}$ is either $(\mathsf{p}_{2},\mathsf{r}_{5})$ or $(\mathsf{p}_{2},\mathsf{r}_{1})$. Note that we are necessarily in the second case since we assumed that the formula holds and since $(\mathsf{p}_{1},\mathsf{q}_{1})(\mathsf{q}_{0},\mathsf{r}_{2})\in\texttt{view}_{3}(p)$. This implies that $(\mathsf{p}_{1},\mathsf{q}_{1})(\mathsf{q}_{2},\mathsf{r}_{0})(\mathsf{r}_{1},\mathsf{p}_{2})\in\texttt{view}_{3}(p)$, and thus $(\mathsf{p}_{1},\mathsf{q}_{1})(\mathsf{q}_{2},\mathsf{r}_{0})(\mathsf{r}_{1},\mathsf{p}_{0})\notin\texttt{view}_{3}(p)$, contradicting the fact that the formula holds.
From Claim 1, $qr$ must be labelled $(\mathsf{q}_{0},\mathsf{r}_{0})$, $(\mathsf{q}_{0},\mathsf{r}_{2})$, $(\mathsf{q}_{2},\mathsf{r}_{0})$ or $(\mathsf{q}_{2},\mathsf{r}_{2})$. Applying Claim 2 to each possibility, $p$ can detect the label of $qr$ and thus infer the chirality of $q$.
3 A Proof Labelling Scheme for Leader Election
Our aim is to orient the edges between particles so that a unique sink particle (i.e., particle with no outgoing edges) that we define to be the leader exists. The certificate given to each particle consists of a direction for each edge incident to the particle. The orientation of the edges is chosen so that particles that are reached by an outgoing edge of some particle $p$, induce a connected graph of size at most 3. In general we cannot avoid the existence of cycles in the orientation, but we will show that the existence of a unique sink is always guaranteed. Each particle $p$ can check that the following rules are locally satisfied or detect an error.
  1. R1
    Each edge is oriented and both particles agree on the direction of the edge.
  2. R2
    Particle $p$ has at most three outgoing edges. We consider edges between $p$ and empty nodes to be incoming for $p$.
  3. R3
    When looking at the ports of $p$ cyclically, all outgoing edges of $p$ are consecutive.
  4. R4
    For every 3-particle triangle $p$ belongs in, the triangle is not a cycle.
We call a configuration where every particle satisfies R1–R4 a valid configuration. Note that R4 does not guarantee an acyclic orientation (i.e., that larger cycles do not exist in the configuration). We do not forbid global cycles, but we will prove that even if cycles of size larger than three are formed by the incoming and outgoing edges, the remaining rules guarantee that there exists a unique sink in the system that we define to be the leader.
Theorem 3.1.
If all rules R1–R4 are satisfied, then there exists a unique sink in the system.
A valid configuration that does not contain any oriented cycle is a valid acyclic orientation. Observe that any particle system admits a valid acyclic orientation, as it can be constructed from any erosion based Leader Election algorithms for programmable matter (e.g., [16, 21]). Indeed, consider an execution of an erosion algorithm on a system, and orient any edge $pq$ from $p$ to $q$ if $p$ is eroded before $q$ in the execution. This orientation is acyclic and it thus obviously satisfies R1 and R4. Since an erosion based algorithm erodes only a particle that does not disconnect its neighbourhood and that is strictly convex (i.e., that has at most three non-eroded neighbours), the orientation also satisfies R2 and R3.
4 A Self-Stabilising Algorithm for Leader Election
In Section 3, we claimed that if all rules are locally satisfied a unique sink exists in the system. Here, we show how a valid configuration is reached from a configuration containing errors. Our algorithm is simple: when a particle, $p$, is incident to an undirected edge $e$, $p$ orients $e$ as outgoing. If the orientation of the edges incident to $p$ violates a rule, $p$ undirects all its outgoing edges. Each activated particle always executes both lines of Algorithm 1.
If $\neg$¬R1 : Mark all undirected edges as outgoing
If $\neg$¬R2 $\lor$ $\neg$¬R3 $\lor$ $\neg$¬R4 : Mark all outgoing edges as undirected
Algorithm 1 Self Stabilising LE
[…]
Theorem 4.1.
Starting from an arbitrary simply connected configuration any Gouda fair execution of Algorithm 1 eventually reaches a configuration satisfying R1–R4 in which no rules can be applied and there exists a unique sink.
5 Proof of Theorem 3.1 and Theorem 4.1
Here, we prove Theorem 4.1. Notice the second statement of Theorem 4.1 is precisely Theorem 3.1. From Remark 2.1, we can consider sequential Gouda fair executions in the remainder of the paper. A configuration of a particle system executing Algorithm 1 is described by the direction of each edge $pq$ (i.e., $\overrightarrow{pq}$, $\overrightarrow{qp}$ or undirected). We make a few observations on how to change the orientation of some edges of a valid configuration and maintain a valid configuration.
{observation}
Let $p$ be a particle such that R3 is satisfied at $p$. Let $e$ be an incoming edge to $p$ and $e^{\prime}$ be an outgoing edge of $p$, s.t. when moving cyclically around $p$, $e$ and $e^{\prime}$ are consecutive. If $e$ becomes outgoing (resp. $e^{\prime}$ becomes incoming), R3 is not violated at $p$.
{observation}
Let $C$ be a configuration and let $p$ be a particle so that R1 (resp., R2, R4) is satisfied at some particle $q\neq p$ in $C$. Then, R1 (resp., R2, R4) is satisfied at $q$ in $C\setminus\{p\}$.
Let $\mathcal{S}=C_{0},C_{1},\ldots$ be an execution of Algorithm 1 starting from a configuration $C_{0}$. Notice a particle $p$ is activable in a configuration $C$ if when it executes Algorithm 1, one of its undirected edges becomes outgoing or one of its outgoing edges becomes undirected. If there exists a configuration $C_{f}$ where no node is activable, then $C_{f}=C_{j}$ for all $j>f$, and we say that the execution stabilises to a final configuration. If all rules are satisfied in this final configuration, then this configuration is valid and we say it is a final directed configuration. If a configuration $C_{i}$ is not final, we can assume that there exists an activable particle $p_{i}$ such that we obtain $C_{i+1}$ by activating $p_{i}$ in $C_{i}$. Since each particle has constant memory, the number of possible configurations is finite. Hence there exists an index $i_{0}$ in $\mathcal{S}$ such that any configuration $C_{i}$ with $i\geq i_{0}$ appears infinitely often in $\mathcal{S}$. We write $\mathcal{S}_{i_{0}}=C_{i_{0}},C_{i_{0}+1},\ldots$ to denote the part of the execution starting at $C_{i_{0}}$ and in the following we consider only $\mathcal{S}_{i_{0}}$ and configurations $C_{i}$ with $i\geq i_{0}$. We call the edges that are never undirected in $\mathcal{S}_{i_{0}}$, stable edges. Observe that by the definition of $i_{0}$, each edge is either stable or undirected infinitely often. Notice that any edge $pq$ directed from $p$ to $q$ in $C_{i}$ with $i\geq i_{0}$, is directed from $p$ to $q$ infinitely often, regardless of whether it is stable.
We first establish a few properties satisfied in $\mathcal{S}_{i_{0}}$.
Lemma 5.1.
In any configuration $C_{i}$ with $i\geq i_{0}$, Rules R2, R3 and R4 are always satisfied.
Proof 5.2.
Let $n_{i}$ be the number of particles in $C_{i}$ that do not satisfy R2, R3 or R4. If a particle $p$ is activated in $C_{i}$, then R2, R3 and R4 are satisfied at $p$ in $C_{i+1}$. Moreover, if R2, R3, and R4 are satisfied at some particle $p$ in $C_{i}$ that is not activated at step $i$, then they are still satisfied at step $C_{i+1}$. Consequently, $n_{i+1}\leq n_{i}$. Since for $i\geq i_{0}$, $C_{i}$ appears infinitely often, we get that for every $i\geq i_{0}$, we have $n_{i}=n_{i_{0}}$. If $n_{i}>0$, there exists a particle $p$ that always violates one of the rules R2, R3 or R4. Thus, $p$ is eventually activated at some step $i$ and in $C_{i+1}$, $p$ satisfies the rules, a contradiction. Consequently, for any $i\geq i_{0}$, R2, R3 and R4 are satisfied at every particle in $C_{i}$.
Lemma 5.3.
If a particle $p$ is incident to a stable outgoing edge, $p$ is never activable in $\mathcal{S}_{i_{0}}$ and all edges incident to $p$ are stable edges.
Proof 5.4.
In a configuration $C_{i}$, if a particle $p$ is incident to an outgoing edge and an undirected edge, then $p$ is activable in $C_{i}$. After its activation, either all the undirected edges incident to $p$ have become outgoing edges, or all outgoing edges of $p$ have become undirected.
Let $\overrightarrow{pq}$ be a stable edge, hence, $p$ never marks $\overrightarrow{pq}$ as undirected. Let us suppose that in addition to $\overrightarrow{pq}$, $p$ is also incident to an unstable edge $pr$. Then infinitely often, $pr$ is undirected and thus there exists a step where $p$ is activated and $pr$ is undirected. At this step, $p$ marks $\overrightarrow{pr}$ as outgoing. Then if $\overrightarrow{pr}$ becomes undirected at a later step, $\overrightarrow{pq}$ must also become undirected, which is a contradiction. Hence, all edges incident to $p$ are stable.
Lemma 5.5.
If a particle $p$ is incident to an unstable edge in $\mathcal{S}_{i_{0}}$, the unstable edges incident to $p$ are at least two and do not appear consecutively around $p$, or there are at least four unstable edges incident to $p$.
Proof 5.6.
Suppose the lemma does not hold and that there exists a particle $p$ incident to $1\leq k\leq 3$ unstable edges that appear consecutively around $p$. Then, there exists an unstable edge $pq$ incident to $p$ such that for every unstable edge $pr$ incident to $p$, either $r=q$ or $r$ is adjacent to $q$. Note that by Lemma 5.3, all stable edges incident to $p$ are incoming to $p$.
Claim 3.
Consider an unstable edge $pr$ with $r\neq q$ and let $s$ be the common neighbour of $p$ and $r$ that is distinct from $q$. Then $\overrightarrow{sp}$ and $\overrightarrow{sr}$ are stable.
{claimproof}
By the definition of $q$, ${sp}$ is stable, and by Lemma 5.3, $sp$ is oriented from $s$ to $p$. By Lemma 5.3 applied at $s$ and $r$, $sr$ is also stable and it is directed from $s$ to $r$
Suppose first there is a configuration $C_{i}$ with $i\geq i_{0}$ such that $pq$ is directed from $p$ to $q$ in $C_{i}$ and undirected in $C_{i+1}$. This implies that $p$ is activated at step $i$. By Lemma 5.1 there exists at least an undirected edge $pr$ in $C_{i}$, and when orienting all undirected edges incident to $p$ in $C_{i}$ as outgoing edges, one of R2, R3, R4 is violated. By the definition of $q$, this cannot be R2 or R3. If R4 is violated, it implies that in $C_{i}$, there exists an undirected edge $pr$ and directed edges $\overrightarrow{rs}$ and $\overrightarrow{sp}$. By Claim 3, $s=q$ but this is impossible since $\overrightarrow{pq}$ is in $C_{i}$.
Then, at each step $i\geq i_{0}$, either $qp$ is undirected or it is directed from $q$ to $p$. If there is no step $i\geq i_{0}$ where $q$ is activated, then $q$ never has any outgoing edge, $qp$ is always undirected and we let $i_{1}=i_{0}$. Otherwise, consider a step $i_{1}-1$ where $q$ is activated such that in $C_{i_{1}}$, $qp$ is undirected. Then at step $i_{1}$, all edges incident to $q$ are either incoming or undirected. We claim that if we activate $p$ at step $i_{1}$, it orients $pq$ from $p$ to $q$. Indeed by the definition of $q$, rules R2 and R3 are satisfied when $p$ orients all its outgoing edges. By Claim 3, any triangle violating R4 should contain $q$, but this is impossible since $q$ has no outgoing edges in $C_{i_{1}}$. So, by the fairness condition, there exists a configuration $C_{i}$ containing $\overrightarrow{pq}$, a contradiction.
We now prove Theorem 4.1 using the structure of the boundary of $\mathcal{P}$ given by Lemma 2.3. Informally, the proof has the following structure. We assume that it is possible that the system does not stabilise and we will arrive at a contradiction. Out of the particle systems that do not stabilise to a configuration that satisfies all rules and has a unique sink, we take a particle system with the minimum number of particles. On the boundary of that system there exists a particle $p$ satisfying one of the cases of Lemma 2.3. For each orientation of the edges incident to $p$ we show that the edges incident to $p$ are stable. Then we take a smaller system that contains exactly one less particle, $p$. We show that the execution in both systems for particles that are not $p$ is the same and as a result, if the system that contains $p$ does not satisfy all rules and does not have a unique sink, then the same is true for the system that does not contain $p$. Since we had assumed that the system containing $p$ is the minimum size system that does not stabilise to a valid configuration, a smaller system not stabilising is a contradiction.
Proof 5.7 (Proof of Theorem 4.1).
Let us suppose that there exists a fair execution $\mathcal{S}=C_{0},C_{1},\ldots$ on a particle configuration $C=C_{0}$ that does not stabilise to a final directed configuration containing a unique sink. Consider such an execution $\mathcal{S}$ with a support $\mathcal{P}$ of minimum size. As defined above, consider a fair execution $\mathcal{S}_{i_{0}}=C_{i_{0}},C_{i_{0}+1},\ldots$ containing only configurations appearing infinitely often. By Lemma 2.3, we can assume that the boundary of $\mathcal{P}$ contains either a pending particle, or a $60^{\circ}$ particle, or two $120^{\circ}$ particles that are connected by a path of $180^{\circ}$ particles on the boundary. In the following, we show that each of these cases cannot occur. We first consider the case where $\mathcal{P}$ contains a pending particle.
Lemma 5.8.
If $\mathcal{P}$ contains a pending particle $p$ (i.e., a particle with only one neighbouring particle $w$), all edges are stable in $\mathcal{S}_{i_{0}}$ and there is a unique sink in the final configuration.
Proof 5.9.
By Lemma 5.5, the edge $pw$ is stable. Suppose first that $pw$ is directed from $p$ to $w$ in $\mathcal{S}_{i_{0}}$. For each $i\geq i_{0}$, let $C^{\prime}_{i}=C_{i}\setminus\{p\}$ and consider the sequence of configurations $\mathcal{S}^{\prime}_{i_{0}}=C^{\prime}_{i_{0}},C^{\prime}_{i_{0}+1},\ldots,C^{\prime}_{i},\ldots$. Observe that for each $i\geq i_{0}$, either $C_{i+1}=C_{i}$ or there exists $p_{i}$ such that $C_{i+1}$ is obtained from $C_{i}$ by activating $p_{i}$ and thus modifying the orientations of edges incident to $p_{i}$. Since $\overrightarrow{pw}$ is stable, by Lemma 5.3, for any $i\geq i_{0}$, $p_{i}\neq p$. Moreover, for each $i\geq i_{0}$, the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. So, $p^{\prime}\neq p$ is activable in $C_{i}^{\prime}$ if and only if it is activable in $C_{i}$. Furthermore, the configuration obtained by activating $p_{i}$ in $C_{i}^{\prime}$ is precisely $C_{i+1}^{\prime}$ since the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. Hence, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}$. Since the edges incident to $p$ are stable, $C_{i_{1}}$ is a final configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $p$ has an outgoing edge, $\overrightarrow{pw}$, in $C_{i_{1}}=C_{i_{0}}$, $p$ is not a sink of $C_{i_{0}}$ and $p^{\prime\prime}$ is the unique sink in $C_{i_{0}}$.
Suppose now that $\overrightarrow{wp}$ is stable. Notice that in this case, $p$ is a sink in $C_{i}$ for each $i\geq i_{0}$. Moreover, since $\overrightarrow{wp}$ is stable, by Lemma 5.3, $w$ is never activated. Since the two common neighbours of $p$ and $w$ are empty, by R3, $p$ is the only outgoing neighbour of $w$ in $C_{i}$ for any $i\geq i_{0}$. Consequently, $w$ is a sink in $C_{i}\setminus\{p\}$ for any $i\geq i_{0}$. For each $i\geq i_{0}$, let $C^{\prime}_{i}=C_{i}\setminus\{p\}$ and consider the sequence of configurations $\mathcal{S}^{\prime}_{i_{0}}=C^{\prime}_{i_{0}},C^{\prime}_{i_{0}+1},\ldots,C^{\prime}_{i},\ldots$. Observe that for each $i\geq i_{0}$, either $C_{i+1}=C_{i}$ or there exists $p_{i}$ such that $C_{i+1}$ is obtained from $C_{i}$ by activating $p_{i}$ and thus modifying the orientations of edges incident to $p_{i}$. Since $p$ has only incoming edges in $C_{i}$, $p_{i}\neq p$. Moreover, since $\overrightarrow{wp}$ is stable, $p_{i}\neq w$. Furthermore, for each $i\geq i_{0}$, the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. Consequently, $p^{\prime}\neq p$ is activable in $C_{i}^{\prime}$ if and only if it is activable in $C_{i}$. Furthermore, the configuration obtained by activating $p_{i}$ in $C_{i}^{\prime}$ is precisely $C_{i+1}^{\prime}$ since the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. Consequently, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}=w$ in $C_{i_{1}}^{\prime}$. Since $\overrightarrow{wp}$ is stable, $C_{i_{1}}$ is a final configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since any $p^{\prime}\in C_{i_{0}}\setminus\{p,w\}$ is not a sink in $C^{\prime}_{i_{0}}$, and since $\overrightarrow{wp}$ is in $C_{i_{0}}$, $p$ is the unique sink in the valid configuration $C_{i_{0}}$.
We now consider the case where the boundary of $\mathcal{P}$ contains a $60^{\circ}$ particle $p$, and we let $q$ and $r$ be the two neighbours of $p$ on the boundary of $\mathcal{P}$.
Lemma 5.10.
If $\mathcal{P}$ contains a $60^{\circ}$ particle $p$, all edges are stable and there is a unique sink in the final configuration.
Proof 5.11.
By Lemma 5.5, $pq$ and $pr$ are stable. Consequently, $pq$ and $pr$ are always directed in the same way all along $\mathcal{S}_{i_{0}}$ and we can talk about the orientation of $pq$ and $pr$ in $\mathcal{S}_{i_{0}}$. For each $i\geq i_{0}$, let $C^{\prime}_{i}=C_{i}\setminus\{p\}$ and consider the sequence of configurations $\mathcal{S}_{i_{0}}^{\prime}=C^{\prime}_{i_{0}},C^{\prime}_{i_{0}+1},\ldots,C^{\prime}_{i},\ldots$. For each $i\geq i_{0}$, either $C_{i+1}=C_{i}$ or there exists $p_{i}$ such that $C_{i+1}$ is obtained from $C_{i}$ by activating $p_{i}$ and thus modifying the orientations of edges incident to $p_{i}$. Since all edges incident to $p$ are stable in $\mathcal{S}_{i_{0}}$, we can assume $p_{i}\neq p$, for any $i\geq i_{0}$.
We distinguish three cases, depending on the orientation of $pq$ and $pr$ in $\mathcal{S}_{i_{0}}$.
Case 1.
The edges incident to $p$ are $\overrightarrow{pq}$ and $\overrightarrow{pr}$.
{claimproof}
For each $i\geq i_{0}$, the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. Hence, a particle $p^{\prime}\neq p$ is activable in $C_{i}^{\prime}$ if and only if it is activable in $C_{i}$. The configuration obtained by activating $p_{i}$ in $C_{i}^{\prime}$ is precisely $C_{i+1}^{\prime}$ since the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. So, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration with a unique sink $p^{\prime\prime}$ where all edges are oriented. Since the edges incident to $p$ are stable, $C_{i_{1}}$ is a final configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $p$ has only outgoing edges in $C_{i_{1}}=C_{i_{0}}$, $p$ is not a sink of $C_{i_{0}}$ and $p^{\prime\prime}$ is the unique sink in $C_{i_{0}}$.
Case 2.
The edges incident to $p$ are $\overrightarrow{qp}$ and $\overrightarrow{pr}$.
{claimproof}
Since $\overrightarrow{qp}$ is stable, $qr$ is also stable by Lemma 5.3. By R4, $qr$ is directed from $q$ to $r$ in $\mathcal{S}_{i_{0}}$. Notice that since $q$ and $p$ are incident to outgoing stable edges, from Lemma 5.3, $p$ and $q$ are incident only to stable edges and are never activable. The edges of $C^{\prime}_{i\geq i_{0}}\setminus\{p\}$ have the same orientation as in $C_{i\geq i_{0}}$. Consequently, for any $p^{\prime}\notin\{p,q\}$, $p^{\prime}$ is activable in $C^{\prime}_{i}$ if and only if it is activable in $C_{i}$. Let us consider $q$. In $C_{i}$, $\overrightarrow{qp}$ is stable and directed and in $C^{\prime}_{i}$, $q$ has an incoming edge from the respective empty node. Furthermore, $q$ is incident to an incoming edge from the empty common neighbour of $p$ and $q$. Hence, R3 is satisfied for $q$ in $C^{\prime}_{i}$ from Observation 5 and the remaining rules are satisfied for $q$ in $C^{\prime}_{i}$ from Observation 5. Therefore, $q$ is never activable in $C^{\prime}_{i\geq i_{0}}$ and thus a particle $p^{\prime}\neq p$ is activable in $C^{\prime}_{i\geq i_{0}}$ if and only if it is activable in $C_{i\geq i_{0}}$. Consequently, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a stable configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}$. Note that $p^{\prime\prime}\neq q$, since $\overrightarrow{qr}\in C_{i_{1}}^{\prime}$. Since the edges incident to $p$ are stable and since $q$ is not activable, $C_{i_{1}}$ is a stable configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $p$ has an outgoing edge in $C_{i_{1}}=C_{i_{0}}$, $p$ is not a sink of $C_{i_{0}}$ and $p^{\prime\prime}$ is therefore the unique sink in $C_{i_{0}}$.
Case 3.
The edges incident to $p$ are $\overrightarrow{qp}$ and $\overrightarrow{rp}$.
{claimproof}
Since $\overrightarrow{qp}$ and $\overrightarrow{rp}$ are stable, from Lemma  5.3, $q$ and $r$ are never activable and all edges incident to $q$ and $r$ are stable. Hence $qr$ is stable and we assume without loss of generality that $qr$ is directed as $\overrightarrow{qr}$ in $C_{i\geq i_{0}}$.
Notice that $p$ is a sink in $C$ and that from R3 all edges incident to $r$ except $\overrightarrow{rp}$ are incoming to $r$ in $C_{i}$. Edges in $C^{\prime}_{i}\setminus\{p\}$ have the same orientation as in $C_{i}$. Using the arguments in Case 2, R1–R4 are satisfied for $q$ and for $r$ in $C^{\prime}_{i}$. So, $q$ and $r$ are never activable in $C^{\prime}_{i}$. For the reasons in Case 2, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. Since $\mathcal{P}$ is of minimum size, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a stable configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}$. Since the only outgoing edge of $r$ in $C_{i_{0}}$ is $\overrightarrow{rp}$, $r=p^{\prime\prime}$ is the unique sink of $C_{i_{1}}^{\prime}$. Furthermore, since the edges incident to $p,q,r$ are stable, $C_{i_{1}}$ is a final configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $p$ has only incoming edges in $C_{i_{1}}=C_{i_{0}}$, $p$ is a sink in $C_{i_{0}}$ and $r$ is not a sink in $C_{i_{0}}$ due to $\overrightarrow{rp}$, so $p$ is the unique sink in $C_{i_{0}}$. Therefore, for any orientation of the edges incident to $p$ in $\mathcal{S}_{i_{0}}$, $C_{i_{0}}$ is a directed final configuration containing a unique sink.
Now assume there exist two $120^{\circ}$ particles connected by a path of $180^{\circ}$ particles on the boundary of $\mathcal{P}$.
Lemma 5.12.
If $\mathcal{P}$ contains two $120^{\circ}$ particles $p,p^{*}$ connected by a path of $180^{\circ}$ particles on the boundary, all edges are stable and there is a unique sink in the final configuration.
Proof 5.13.
Let $q$ and $r$ be the neighbours of $p$ on the boundary and let $s$ be the common neighbour of $p,q$ and $r$. By Lemma 5.5, we know that $ps$ is stable in $\mathcal{S}_{i_{0}}$. We split the proof of the lemma in different cases, depending on the orientation of $ps$ in $C_{i\geq i_{0}}$.
Case 1.
The edge between $p$ and $s$ is $\overrightarrow{ps}$.
For each $i\geq i_{0}$, let $C^{\prime}_{i}=C_{i}\setminus\{p\}$ and consider the sequence of configurations $\mathcal{S}_{i_{0}}^{\prime}=C^{\prime}_{i_{0}},C^{\prime}_{i_{0}+1},\ldots,C^{\prime}_{i},\ldots$ For each $i\geq i_{0}$, either $C_{i+1}=C_{i}$ or there exists $p_{i}$ such that $C_{i+1}$ is obtained from $C_{i}$ by activating $p_{i}$ and thus modifying the orientations of edges incident to $p_{i}$. From Lemma 5.3, since $\overrightarrow{ps}$ is stable, $p$ is never activable and the edges $pq$ and $pr$ are stable. Consequently, $p_{i}\neq p$ for any $i\geq i_{0}$. The orientations of $pq$ and $pr$ lead to the following cases.
Case 1.1.
Particle $p$ is incident to $\overrightarrow{pq}$, $\overrightarrow{ps}$ and $\overrightarrow{pr}$.
{claimproof}
For each $i\geq i_{0}$, the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. So, a particle $p^{\prime}\neq p$ is activable in $C_{i}^{\prime}$ if and only if it is activable in $C_{i}$. The configuration obtained by activating $p_{i}$ in $C_{i}^{\prime}$ is precisely $C_{i+1}^{\prime}$ since the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. Hence, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration with a unique sink $p^{\prime\prime}$ where all edges are oriented. Since the edges incident to $p$ are stable, $C_{i_{1}}$ is a final configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $p$ has only outgoing edges in $C_{i_{1}}=C_{i_{0}}$, $p$ is not a sink of $C_{i_{0}}$ and $p^{\prime\prime}$ is the unique sink in $C_{i_{0}}$.
Case 1.2.
Particle $p$ is incident to $\overrightarrow{pq}$, $\overrightarrow{ps}$ and $\overrightarrow{rp}$.
{claimproof}
Since $\overrightarrow{rp}$ is stable, from Lemma 5.3, $r$ is not activable in $C$ and $rs$ is stable. From R4, $\overrightarrow{rs}$ is in $C_{i\geq i_{0}}$. The edges of $C^{\prime}_{i\geq i_{0}}\setminus\{p\}$ have the same orientation as in $C_{i\geq i_{0}}$. So for any $p^{\prime}\notin\{p,r\}$, $p^{\prime}$ is activable in $C_{i}^{\prime}$ if and only if $p^{\prime}$ is activable in $C_{i}$. Let us consider $r$. From Observation 5 and the incoming edge from the empty common neighbour of $p$ and $r$ to $r$, R3 is satisfied for $r$ in $C^{\prime}_{i}$. R1, R2 and R4 are satisfied for $r$ in $C^{\prime}_{i}$ from Observation 5. Hence, $r$ is never activable in $C^{\prime}_{i}$. Furthermore, the configuration obtained by activating $p_{i}\neq r$ in $C_{i}^{\prime}$ is precisely $C_{i+1}^{\prime}$ since the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. So, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}\neq r$, since $r$ has an outgoing edge $\overrightarrow{rs}$ in $C^{\prime}_{i}$. Since the edges incident to $p$ are stable, $C_{i_{1}}$ is a stable configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $p$ has outgoing edges in $C_{i_{1}}=C_{i_{0}}$, $p$ is not a sink of $C_{i_{0}}$ and $p^{\prime\prime}$ is the unique sink in $C_{i_{0}}$.
Case 1.3.
Particle $p$ is incident to $\overrightarrow{qp}$, $\overrightarrow{ps}$ and $\overrightarrow{rp}$.
{claimproof}
Since $\overrightarrow{qp}$ and $\overrightarrow{rp}$ are stable, from Lemma 5.3, $r$ and $q$ are not activable in $C$ and the edges $rs$ and $qs$ are stable. By R4, $\overrightarrow{rs}$ and $\overrightarrow{qs}$ belong to $C_{i\geq i_{0}}$. The edges of $C^{\prime}_{i\geq i_{0}}\setminus\{p\}$ have the same orientation as in $C_{i\geq i_{0}}$. Consequently, for any $p^{\prime}\notin\{p,q,r\}$, $p^{\prime}$ is activable in $C_{i}^{\prime}$ if and only if $p^{\prime}$ is activable in $C_{i}$. Let us consider $q$ and $r$. Using the same arguments as in Case 1.2, R1–R4 are satisfied for $q$ and for $r$ in $C^{\prime}_{i}$. Consequently, $q$ and $r$ are never activable in $C^{\prime}_{i}$. Furthermore, the configuration obtained by activating $p_{i}\notin\{p,q,r\}$ in $C_{i}^{\prime}$ is precisely $C_{i+1}^{\prime}$ since the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. Therefore, $C^{\prime}_{i}$ satisfies all rules and $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}$. Since $q$ and $r$ have both an outgoing edge to $s$ in $C^{\prime}_{i}$, $p^{\prime\prime}\notin\{q,r\}$. Since the edges incident to $p$ are stable, $C_{i_{1}}$ is a final configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $p$ has outgoing edges in $C_{i_{1}}=C_{i_{0}}$, $p$ is not a sink of $C_{i_{0}}$ and $p^{\prime\prime}$ is therefore the unique sink in $C_{i_{0}}$.
We now consider the case where the edge between $p$ and $s$ is $\overrightarrow{sp}$. Observe that by Lemma 5.3, $sq$ and $sr$ are stable edges. Note that by Lemma 5.5, $pq$ is stable if and only if $pr$ is stable. We distinguish two cases depending on whether these two edges are stable.
Case 2.
The edge between $p$ and $s$ is $\overrightarrow{sp}$, and the edges $pq$ and $pr$ are stable.
The possible orientations of the stable edges $sq$ and $sr$ in $\mathcal{S}_{i_{0}}$ give the following subcases.
Case 2.1.
Particle $s$ is incident to $\overrightarrow{qs}$ and $\overrightarrow{rs}$.
{claimproof}
By Lemma 5.3, $r$ and $q$ are never activable and $rp$ and $qp$ are stable. By R4, $\overrightarrow{qp}$ and $\overrightarrow{rp}$ are in $C_{i\geq i_{0}}$.
The edges of $C^{\prime}_{i\geq i_{0}}\setminus\{p\}$ have the same orientation as in $C_{i\geq i_{0}}$ and thus for any $p^{\prime}\notin\{q,r,s\}$, $p^{\prime}$ is activable in $C^{\prime}_{i}$ if and only if it is activable in $C_{i}$. Let us consider $q,r,s$. Using the same arguments as in Case 1.2, R1–R4 are satisfied for $q$ and $r$ in $C\setminus\{p\}$. Since $sp$ is between two incoming edges from Observation 5, R3 is satisfied for $s$ in $C^{\prime}_{i}$. R1, R2 and R4 are satisfied for $s$ in $C^{\prime}_{i}$ from Observation 5. So, $q,r,s$ are never activable in $C^{\prime}_{i}$. Furthermore, the configuration obtained by activating $p_{i}\neq q,s,r$ in $C_{i}^{\prime}$ is precisely $C_{i+1}^{\prime}$ since the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. Hence, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration with a unique sink $p^{\prime\prime}$ where all edges are oriented. Notice that by R3, the only outgoing edge of $s$ in $C_{i}$ is $\overrightarrow{sp}$, so $p^{\prime\prime}=s$ is the unique sink of $C^{\prime}_{i}$. Since the edges incident to $p$ are stable, $C_{i_{1}}$ is a final configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $s$ has an outgoing edge $\overrightarrow{sp}$ in $C_{i_{1}}=C_{i_{0}}$, $s$ is not a sink of $C_{i_{0}}$ and $p$ is the unique sink in $C_{i_{0}}$.
Observe that if $s$ is incident to $\overrightarrow{qs}$ (resp., $\overrightarrow{rs}$), then since $qp$ (resp., $rp$) is stable, by R4, $p$ is incident to $\overrightarrow{qp}$ (resp., $\overrightarrow{rp}$). When among $qs$ and $rs$, there is one outgoing and one incoming edge, we consider two cases depending on whether $p$ is a sink in $\mathcal{S}_{i_{0}}$.
Case 2.2.
Particle $s$ is incident to $\overrightarrow{qs}$ and $\overrightarrow{sr}$ and $p$ is incident to $\overrightarrow{rp}$.
{claimproof}
Since $\overrightarrow{qs}$ and $\overrightarrow{sp}$ are stable, necessarily $\overrightarrow{qp}$ is stable and $p$ is a sink in $C_{i\geq i_{0}}$. So, $p$ is never activable in $\mathcal{S}_{i_{0}}$ and from Lemma 5.3, $q,r,s$ are never activable in $\mathcal{S}_{i_{0}}$ either.
The edges of $C^{\prime}_{i\geq i_{0}}\setminus\{p\}$ have the same orientation as in $C_{i\geq i_{0}}$ and thus any particle $p^{\prime}\notin\{p,q,r,s\}$ is activable in $C^{\prime}_{i}$ if and only if it is activable in $C_{i}$. Let us consider $q,r,s$. Since $\overrightarrow{sp}$ is between an incoming and an outgoing edge, by Observation 5, R3 is satisfied for $s$ in $C^{\prime}_{i}$ and the remaining rules are satisfied for $s$ in $C^{\prime}_{i}$ from Observation 5. The arguments for $q$ are the same as in Case 2.1. Due to the incoming edge from the common empty neighbour of $r$ and $p$ and from Observation 5, R3 is satisfied for $r$ in $C^{\prime}_{i}$ and the remaining rules are satisfied for $r$ in $C^{\prime}_{i}$ from Observation 5. So, $q,r,s$ are never activable in $C^{\prime}_{i\geq i_{0}}$. Furthermore, the configuration obtained by activating $p_{i}\notin\{q,r,s\}$ in $C_{i}^{\prime}$ is precisely $C_{i+1}^{\prime}$ since the edges of $C_{i}^{\prime}$ have the same orientation as in $C_{i}$. Hence, all rules are satisfied in $C^{\prime}_{i}$ and $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}$. Notice that by R3, the only outgoing edge of $r$ in $C_{i}$ is $\overrightarrow{rp}$, so $p^{\prime\prime}=r$ is the unique sink of $C^{\prime}_{i}$. As the edges incident to $p$ are stable, $C_{i_{1}}$ is a final configuration where all edges are oriented. By our definition of $i_{0}$, $i_{1}=i_{0}$. Since $p$ does not have outgoing edges in $C_{i_{1}}=C_{i_{0}}$, $r$ is not a sink of $C_{i_{0}}$ due to $\overrightarrow{rp}$ and $p$ is the unique sink in $C_{i_{0}}$.
Case 2.3.
Particle $s$ is incident to $\overrightarrow{qs}$ and $\overrightarrow{sr}$ and $p$ is incident to $\overrightarrow{pr}$.
{claimproof}
As noted before the stable edge $qp$ is oriented as $\overrightarrow{qp}$ by R4. By Lemma 5.3, $p$, $q$ and $s$ are never activable in $\mathcal{S}_{i_{0}}$ and the edges $pr$ and $sr$ are stable. The edges of $C^{\prime}_{i\geq i_{0}}\setminus\{p\}$ have the same orientation as in $C_{i\geq i_{0}}$. So, for any $p^{\prime}\notin\{p,q,s\}$, $p^{\prime}$ is activable in $C^{\prime}_{i}$ if and only if it is activable in $C_{i}$. Let us consider $q$ and $s$. Using the same arguments for both $q$ and $s$ as in Case 2.2, we obtain that R1–R4 are always satisfied at $q$ and $s$, and that they are never activable in $C_{i\geq i_{0}}^{\prime}$. Hence, $\mathcal{S}^{\prime}_{i_{0}}$ is a fair execution of Algorithm 1 on $\mathcal{P}\setminus\{p\}$. By the minimality of the size of $\mathcal{P}$, there exists a step $i_{1}\geq i_{0}$ such that $C_{i_{1}}^{\prime}$ is a final configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}\notin\{q,s\}$, since $q$ and $s$ have outgoing edges in $C\setminus\{p\}$. Since the edges incident to $p$ are stable, $C_{i_{1}}$ is a stable configuration where all edges are oriented. By our definition of $i_{0}$, this implies that $i_{1}=i_{0}$. Since $p$ is incident to $\overrightarrow{pr}$ in $C_{i_{1}}=C_{i_{0}}$, $p$ is not a sink in $C_{i}$, hence $p^{\prime\prime}$ is the unique sink in $C_{i_{0}}$.
Case 2.4.
Particle $s$ is incident to $\overrightarrow{sq}$ and $\overrightarrow{sr}$.
{claimproof}
Since $pq,pr,ps$ are stable and since $ps$ is directed as $\overrightarrow{sp}$ in $C_{i\geq i_{0}}$, $p$ is incident to at most one outgoing edge. Without loss of generality, we can thus assume that $\overrightarrow{qp}$ is in $C_{i\geq i_{0}}$. Observe that by R3, $\overrightarrow{qp}$ is the only outgoing edge at $q$ and no neighbour of $q$ is activable in $C_{i\geq i_{0}}$ by Lemma 5.3. Note also that $s$ has three outgoing edges $\overrightarrow{sq},\overrightarrow{sp},\overrightarrow{sr}$ in $C_{i}$ and since $s$ is not activable in $C_{i}$, by R2, all other edges incident to $s$ are incoming. Again, this implies that all neighbours of $s$ different from $r$ are not activable in $C_{i\geq i_{0}}$. For each $i\geq i_{0}$, let $C^{*}_{i}$ be the configuration obtained from $C_{i}$ by replacing $\overrightarrow{sq}$ by $\overrightarrow{qs}$.
For any particle $p^{\prime}\notin\{s,q\}$, R2 and R3 are satisfied at $p^{\prime}$ in $C^{*}_{i}$ since they are satisfied at $p^{\prime}$ in $C_{i}$ by Lemma 5.1 and the orientation of the edges incident to $p^{\prime}$ in $C_{i}^{*}$ is the same as in $C_{i}$. Since $q$ only has one outgoing edge in $C_{i}$, R2 and R3 are satisfied at $q$ in $C^{*}_{i}$. Since $s$ has three outgoing edges $\overrightarrow{sq},\overrightarrow{sp},\overrightarrow{sr}$ in $C_{i}$ and since all other edges incident to $s$ are incoming, R2 directly holds at $s$ in $C_{i}^{*}$ and R3 holds at $s$ in $C_{i}^{*}$ since $p,q$ and $r$ are reached through consecutive ports of $s$ by definition. If R4 is not satisfied at some particle $p^{\prime}$ in $C_{i}^{*}$, there is a directed triangle made of the edges $\overrightarrow{qs},\overrightarrow{sp^{\prime}},\overrightarrow{p^{\prime}q}$ in $C_{i}^{*}$. Since the only out-neighbours of $s$ in $C_{i}^{*}$ are $p$ and $r$, necessarily, $p^{\prime}=p$, but this is impossible since $pq$ is oriented from $q$ to $p$. So, R2, R3, R4 are always satisfied in $C^{*}_{i\geq i_{0}}$. Since all edges incident to $q$ and $s$ are stable in $C_{i}$, R1 is also satisfied at $s$ and $q$ in $C_{i}$ and in $C_{i}^{*}$. Hence, $q$ and $s$ are never activable in $C^{*}_{i\geq i_{0}}$. For any $p^{\prime}\notin\{q,s\}$, $p^{\prime}$ is activable in $C_{i}^{*}$ if and only if it is activable in $C_{i}$. Therefore, $\mathcal{S}_{i}^{*}$ is a fair execution of Algorithm 1 on $\mathcal{P}$. Note that when considering $C^{*}_{i\geq i_{0}}$, we are in Case 2.2 or 2.3. So, we know that $C^{*}_{i_{0}}$ is a final configuration where all edges are oriented and that contains a unique sink $p^{\prime\prime}$ different from $q$ and $s$. Since a particle $p^{\prime}$ is activable in $C_{i_{0}}$ if and only if it is activable in $C^{*}_{i_{0}}$, $C_{i_{0}}$ is also a valid configuration, and $p^{\prime\prime}$ is the unique sink of $C_{i_{0}}$.
Finally, we consider the case where the edges $pq$ and $pr$ are not stable. We remind the reader that from Lemma 5.5, $pq$ and $pr$ are either both stable or both unstable.
Case 3.
The edge between $p$ and $s$ is $\overrightarrow{sp}$ and the edges $pq$ and $pr$ are not stable.
{claimproof}
We will prove that this case is not possible.
[…]
Figure 3: Left: The setting in Case 3, that is, a $120^{\circ}$ particle $p$ (square) with $\{pq,pr\}$ unstable and $s$ incident to the directed edges $sq$, $sp$ and $sr$. Right: The final orientation of edges in Case 3.
Without loss of generality, assume that $r$ is on the path connecting $p$ to $p^{*}$ via $180^{\circ}$ particles. Note that it is possible that $p^{*}=r$. Let $(r_{0}=p,r_{1}=r,r_{2},\ldots,r_{k}=p^{*})$ be the path on the boundary from $p$ to $p^{*}$ whose inner particles are all $180^{\circ}$ particles (if $r=p^{*}$, then $k=1$). Let $s_{j+1}$ be the common neighbour of any pair of consecutive particles $r_{j}$ and $r_{j+1}$ with $10\leq j\leq k-$ 1, and observe that $s_{1}=s$. Let $s_{0}=q$ and let $s_{k+1}$ be the neighbour of $r_{k}$ on the boundary that is distinct from $r_{k-1}$. This setting is also shown in Figure 3(a).
Since $s$ is incident to a stable edge, $\overrightarrow{sp}$, from Lemma 5.3 all edges incident to $s$ are stable. By Lemma 5.3 applied at $q$ and $r$ and since $pq$ and $pr$ are not stable, necessarily $\overrightarrow{sq}$ and $\overrightarrow{sr}$ are stable in $\mathcal{S}_{i_{0}}$. Since in this setting $s$ has three outgoing edges, $sq$, $sp$ and $sr$, from R2 $ss_{2}$ is incoming to $s$. So, $s_{2}$ is incident to the stable outgoing edge $\overrightarrow{s_{2}s}$. From Lemma 5.3, $s_{2}$ is not activable and all edges incident to $s_{2}$ are stable and directed. From R4, $s_{2}r$ is oriented from $s_{2}$ to $r$. If $r=p^{*}$ (i.e., if $k=1$), $r$ is incident to only one edge that is not stable, which is impossible from Lemma 5.5. Hence, $k\geq 2$ and $rr_{2}$ is not stable. As $s_{2}$ is not activable, from Lemma 5.3, $s_{2}r_{2}$ is stable. If $s_{2}r_{2}$ is directed from $r_{2}$ to $s_{2}$, $r_{2}$ has a stable outgoing edge and from Lemma 5.3, $rr_{2}$ is stable, which is a contradiction for $r$. So $s_{2}r_{2}$ is directed from $s_{2}$ to $r_{2}$. Generalising, for $i\geq 1$ each particle $s_{i}$ is incident to the stable edges $\overrightarrow{s_{i}s_{i-1}}$, $\overrightarrow{s_{i}r_{i-1}}$ and $\overrightarrow{s_{i}r_{i}}$ and the edge $r_{i-1}r_{i}$ is not stable. Then, for $i=k$ the edge $r_{k-1}r_{k}$ should be the only unstable edge incident to $r_{k}$ which is impossible from Lemma 5.5, a contradiction.
This ends the proof of Lemma 5.12.
The proof of Theorem 4.1 follows from Lemmas 2.3, 5.8, 5.10 and 5.12.
6 Further Remarks
A very nice property of our algorithm is that it is very simple. However, we showed that our algorithm works assuming that the scheduler is Gouda fair. The execution presented in Figure 4 shows that if we consider a sequential unfair scheduler (i.e., we only ask that the scheduler activates an activable particle at each step), there exist periodic executions that never reach a valid configuration. It would thus be interesting to understand if we can design a self-stabilising leader election algorithm for simply connected configurations that is correct even with an unfair scheduler. In the case where this is possible, we believe that this would lead to a much more complex algorithm than our algorithm.
Figure 4: A periodic unfair execution of our algorithm. At each step, the red vertex is activated, and it modifies the status of its incident red edges (i.e., the ones that are not incoming).
Our algorithm heavily uses the geometry of the system and relies on the fact that the support is simply connected. If we suppose that the particles agree on the orientation of the grid, the impossibility results of Dolev et al. [19] no longer hold. One can thus wonder if, assuming that the particles agree on the orientation of the grid, it is possible to design a silent self-stabilising leader election algorithm using constant memory for arbitrary connected configurations. Again, one should use the geometry of the grid in order to overcome the impossibility results of Dolev et al., but it seems to be very challenging.